# Robust Weight Imprinting: Insights from Neural Collapse and Proxy-Based Aggregation

### v1

@@ Abstract
The capacity of a foundation model allows for adaptation to new downstream tasks. Weight imprinting is a universal and efficient method to fulfill this purpose. It has been reinvented several times, but it has not been systematically studied. In this paper, we propose a framework for imprinting, identifying three main components: generation, normalization, and aggregation. This allows us to conduct an in-depth analysis of imprinting and a comparison of the existing work. We reveal the benefits of representing novel data with multiple proxies in the generation step and show the importance of proper normalization. We determine those proxies through clustering and propose a novel variant of imprinting that outperforms previous work. We motivate this by the neural collapse phenomenon – an important connection that we can draw for the first time. Our results show an increase of up to 4% in challenging scenarios with complex data distributions for new classes.
@@ 1 Introduction
We present a unifying framework that enables a systematic comparison of existing imprinting techniques. More precisely, we can generalize prior work by decomposing imprinting into three principal steps (see fig. 1). During generation (GEN) of weights, the network selects representative data samples and generates one or more weight vectors per class (multiple proxies). Normalization (NORM) is crucial, as the network needs to balance its generated weight vectors. Aggregation (AGG) entails the computation of the final output, e.g., a class label.
The efficiency of imprinting enables us to perform a comprehensive analysis with a large number of experiments. We present a new, best-performing imprinting strategy using multi-modal weight imprinting in combination with the correct way of normalization.
@@
Figure 1: Overview of our imprinting framework. The foundation model FM is frozen and shows neural collapse. The weight generator (GEN) uses training data from a novel task T to generate one or more weight vectors (proxies) per class $1,\dotsc,c$. The final output of the test data in T is computed by an aggregation (AGG) mechanism. Embeddings and generated weights are normalized according to NORMpre and NORMpost, respectively. During inference, embeddings are normalized according to NORMinf (not shown here).
@@ 4 Measurement of Neural Collapse
Two important characteristics of NC are variability collapse, i.e., the within-class variability of the penultimate-layer embeddings collapses to zero, and convergence to nearest-mean-classification. We focus on variability collapse ($\mathcal{NC}_{1}$), using the equation 1 from [13]:
@@ 5 Experimental Setup
We use resnet18 [34] and vit_b_16 [35] as FMs, one CNN-based and one Transformer-based architecture. In neural collapse investigations (section 6.3), we also use resnet50 [34] and swin_b [36]. All of them are pre-trained on ImageNet-1K (ILSVRC 2012) [37]. To generate the embeddings, we use PyTorch’s torchvision models.
@@
To find out the best imprinting strategy within our framework, we focus on tasks T created from the datasets MNIST [38], FashionMNIST [39], and CIFAR-10  [^2] https://www.cs.toronto.edu/~kriz/cifar.html, each containing 10 classes. We mainly focus on the three T containing all ten classes. Furthermore, we look at smaller tasks only containing classes $\{0,1,2\}$, and the two tasks containing classes $\{1,3,5,7,9\}$ resp. $\{0,2,4,6,8\}$. This random selection of $3\cdot 4=12$ tasks adds variation to our evaluation.
In the investigations of neural collapse (NC), we also look at the FMs’ pre-training data (ImageNet). As its test set is not available, we use its validation set in $\mathcal{NC}_{1}$ computations. For ImageNet, we relabel data by combining multiple classes into one label to simulate multi-modal class distributions for an in-depth NC analysis. These tasks are called "$d$ in $1$", $d=1,\dotsc,10$, each containing $10$ different labels. More precisely, we take 100 random classes from ImageNet and sequentially map the first $d$ to label $1$, the second $d$ to label $2$, etc., until we reach 10 distinct labels. See fig. 3 for a simplified example. We do this random sampling 10 times which results in $10\cdot 10=100$ tasks.
@@
In total, we run approximately 350 000 experiments, varying across the imprinting components, foundation models, tasks, and seeds. This is feasible with minimal effort as imprinting is a highly efficient method that operates without relying on gradient descent or other non-linear optimization techniques.
@@
Throughout our experiments, the median accuracy on the test set for three different seeds is reported, if not otherwise specified. In sections 6.1 and 6.2, we investigate the imprinting performance by varying the FM (2) and T (12). We then sort the combinations by their final accuracy. There are $2\cdot 12=24$ potentially different ranks for each of the combinations. We show the average rank, average accuracy, and statistical significance in ranking (dis-)agreements through critical difference (CD) diagrams as presented in [40]. [^3] Code used to generate these diagrams can be found here [41]. In the CD diagrams, a thick horizontal line indicates a group of combinations that are not significantly different from each other in terms of accuracy. We consider differences significant if $p<0.05$.
@@ 6.1 Best Imprinting Strategy
Paper NORMpre GEN NORMpost NORMinf AGG Avg. acc. %
Qi et al. [4] L2 mean L2 L2 max 87.72
Hosoda et al. [14] none mean quantile none max 82.15
Janson et al. [22] none mean none none 1-nn 87.65
Ours L2 k-means L2 L2 max 91.47
[…]
[table]A table beside a figure
@@ 6.3 Neural Collapse and Number of Proxies
Figure 11: Average $\mathcal{NC}_{1}$ and $95\%$ confidence intervals of ten random ImageNet label remappings ("$d$ in $1$") for every $d=1,\dotsc,10$. The $\mathcal{NC}_{1}$ for the tasks containing all of MNIST, FashionMNIST, and CIFAR-10 at once are depicted in dots. When compared to the $\mathcal{NC}_{1}$ on the pre-training set (ImageNet), these sets are less collapsed on the Transformer-based architectures. A clear linear relationship across $d$ can be inferred for all FMs, i.e., increased multi-modality implies less collapse.
@@ A.1 Additional Tables and Critical Difference Diagrams
L2 L2 quantile 80.71
none none quantile 80.72
L2 none quantile 80.72
L2 L2 none 84.03
@@
L2 L2 L2 91.47
L2 none L2 91.45
none none L2 91.45
L2 L2 quantile 90.91
L2 L2 none 89.97
L2 none quantile 79.84
none none quantile 79.84
L2 none none 79.24
none none none 79.24

### v2

@@ Abstract
The capacity of a foundation model allows for adaptation to new downstream tasks. Weight imprinting is a universal and efficient method to fulfill this purpose. It has been reinvented several times, but it has not been systematically studied. In this paper, we propose a framework for imprinting, identifying three main components: generation, normalization, and aggregation. This allows us to conduct an in-depth analysis of imprinting and a comparison of the existing work. We reveal the benefits of representing novel data with multiple proxies in the generation step and show the importance of proper normalization. We determine proxies through clustering and propose a novel variant of imprinting that outperforms previous work. We motivate this by the neural collapse phenomenon – an important connection that we can draw for the first time. Our results show an increase of up to 4% in challenging scenarios with complex data distributions for new classes. Finally, we publicly release our code at https://github.com/DATEXIS/multi-imprinting/.
@@ 1 Introduction
We present a unifying framework that enables a systematic comparison of existing imprinting techniques. More precisely, we generalize prior work by decomposing imprinting into three principal steps (see fig. 1). During generation (GEN) of weights, the network selects representative data samples and generates one or more weight vectors per class (proxies). Normalization (NORM) is crucial, as the network needs to balance its generated weight vectors. Aggregation (AGG) entails the computation of the final output, e.g., a class label.
The efficiency of imprinting enables us to perform a comprehensive analysis with a large number of experiments. We present a new, best-performing imprinting strategy using multi-modal weight imprinting in combination with the correct way of normalization, outperforming previously studied methods, as depicted in fig. 2.
@@
Figure 1: Overview of our imprinting framework. The foundation model FM is frozen and shows neural collapse. The weight generator (GEN) uses training data from a novel task T to generate one or more weight vectors (proxies) per class $1,\dotsc,c$. The final output for the test data in T is computed by an aggregation (AGG) mechanism. Embeddings and generated weights are normalized according to NORMpre and NORMpost, respectively. During inference, embeddings are normalized according to NORMinf (not shown here).
Paper NORMpre GEN NORMpost NORMinf AGG Avg. acc. %
Qi et al. [4] L2 mean L2 L2 max 87.72
Hosoda et al. [14] none mean quantile none max 80.71
Janson et al. [15] none mean none none 1-nn 87.64
Ours L2 k-means L2 L2 max 91.48
[…]
[table]A table beside a figure
@@ 4 Measurement of Neural Collapse
Two important characteristics of NC are variability collapse, i.e., the within-class variability of the penultimate-layer embeddings collapses to zero, and convergence to nearest-mean-classification. We focus on variability collapse ($\mathcal{NC}_{1}$) as in [13]:
@@ 5 Experimental Setup
We use resnet18 [34] and vit_b_16 [35] as FMs, one CNN-based and one Transformer-based architecture. In neural collapse investigations (section 6.3), we also work with resnet50 [34] and swin_b [36]. All four models are pre-trained on ImageNet-1K (ILSVRC 2012) [37]. To generate the embeddings, we use PyTorch’s torchvision models.
@@
To find out the best imprinting strategy within our framework, we focus on tasks T created from the datasets MNIST [38], FashionMNIST [39], and CIFAR-10 [40], each containing 10 classes. We mainly focus on the three T containing all ten classes. Furthermore, we look at smaller tasks only containing classes $\{0,1,2\}$, and the two tasks containing classes $\{1,3,5,7,9\}$ resp. $\{0,2,4,6,8\}$. This random selection of $3\cdot 4=12$ tasks adds variation to our evaluations.
In the investigations of neural collapse (NC), we also look at the FMs’ pre-training data (ImageNet). As its test set is not available, we use its validation set in $\mathcal{NC}_{1}$ computations. For ImageNet, we relabel data by combining multiple classes into one label to simulate multi-modal class distributions for an in-depth NC analysis. These tasks are called "$d$ in $1$", $d=1,\dotsc,10$, each containing $10$ different labels. More precisely, we take 100 random classes from ImageNet and sequentially map the first $d$ to label $1$, the second $d$ to label $2$, etc., until we reach 10 distinct labels. See fig. 4 for a simplified illustration. We do this random sampling 10 times which results in $10\cdot 10=100$ tasks. Furthermore, to ensure scale invariance in covariance-based NC measurements, all embeddings are $L^{2}$-normalized before computing $\mathcal{NC}_{1}$.
@@
In total, we run approximately 150 000 experiments, varying across the imprinting components, foundation models, tasks, and seeds. This is feasible with minimal effort as imprinting is a highly efficient method that operates without relying on gradient descent or other non-linear optimization techniques.
@@
Throughout our experiments, the median accuracy on the test set for three different seeds is reported, if not otherwise specified. In sections 6.1 and 6.2, we investigate the imprinting performance by varying the FM (2) and T (12). We then sort the combinations by their final accuracy. There are $2\cdot 12=24$ potentially different ranks for each of the combinations. We show the average rank, average accuracy, and statistical significance in ranking (dis-)agreements through critical difference (CD) diagrams as presented in [41]. [^2] The code used to generate these diagrams is inspired by [42]. In the CD diagrams, a thick horizontal line indicates a group of combinations that are not significantly different from each other in terms of accuracy. We consider differences significant if $p<0.05$.
@@ 6.3 Neural Collapse and Number of Proxies
Figure 11: Average $\mathcal{NC}_{1}$ and $95\%$ confidence intervals of ten random ImageNet label remappings ("$d$ in $1$") for every $d=1,\dotsc,10$. The $\mathcal{NC}_{1}$ for the tasks containing all of MNIST, FashionMNIST, and CIFAR-10 at once are depicted as dots. When compared to $\mathcal{NC}_{1}$ on the pre-training set (ImageNet), these sets are less collapsed on the Transformer-based architectures. A clear linear relationship across $d$ can be inferred for all FMs, i.e., increased multi-modality implies less collapse.
@@ A.1 Additional Tables and Critical Difference Diagrams
L2 L2 quantile 80.70
L2 none quantile 80.71
none none quantile 80.71
L2 L2 none 84.04
@@
L2 none L2 91.44
none none L2 91.44
L2 L2 L2 91.48
L2 L2 quantile 90.86
L2 L2 none 89.99
L2 none quantile 80.36
none none quantile 80.36
L2 none none 79.22
none none none 79.22